OpenStudy (anonymous):

find f '' (x) from : f(x)=(3x -2)^5

OpenStudy (anonymous):

did you find the first derivative?

OpenStudy (anonymous):

so.. how to find the first derivative...?

OpenStudy (anonymous):

Use the chain rule

OpenStudy (anonymous):

f(g(x))' = f'(g(x))g'(x)

OpenStudy (anonymous):

\[f'(x) = 15\, \left( 3\,x-2 \right) ^{4}\]\[f''(x) = 180\, \left( 3\,x-2 \right) ^{3}\]

OpenStudy (anonymous):

maybe you can explain in detail.?

OpenStudy (anonymous):

You differentiate the function twice.

OpenStudy (anonymous):

Are you familiar with the chain rule?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

I hope u can describe for me...

OpenStudy (anonymous):

You should learn a bit about differentiation first before you can work on these problems.

OpenStudy (anonymous):

I have learn about differentiation first may be like this : f (x) = 3x (x ^ 3-1) =3x^4 -3x f'(x)=12x-3

OpenStudy (anonymous):

Think of the chain rule as layes on a cake

OpenStudy (anonymous):

The function you have is a cake

OpenStudy (anonymous):

The first layer is x^5, the second layer is the inner function or 3x-2

OpenStudy (anonymous):

ok... I have started to understand...

OpenStudy (anonymous):

Now, with the chain rule, you differntiate the first layer leaving the inner layer untouched

OpenStudy (anonymous):

Then differntiatie just the inner layer

OpenStudy (anonymous):

fiest u take 3x-2 as y then differentiate y^5 to get 5y^4 then using chain rule differentiate 3x-2 to get 3(5)y^4.. similarly for f''(x)

OpenStudy (anonymous):

\[f(x) = x^5\]\[g(x) = 3x-2\]\[(f(g(x))' = f'(g(x))g'(x)\]\[f'(x) = 5x^4\]\[g'(x) = 3\]\[f'(g(x))g'(x) = 5(3x-2)^4 \cdot 3 = 15 \cdot ( 3x-2)^4\]

OpenStudy (anonymous):

thanks Alchemista...